OpenStudy (anonymous):

PLEASE HELP?!?! x Power Of 2 + 5x + 6 = 0 {2, 3} {-2, -3} {-1, 6} {1, -6}

OpenStudy (maheshmeghwal9):

Do u mean?\[(2+5x+6)^2=0\]

OpenStudy (maheshmeghwal9):

Sorry\[ (2+5x+6)^x=0.\]

OpenStudy (anonymous):

\[X ^{2} + 5x + 6 =0 < No this is what i mean \]

OpenStudy (maheshmeghwal9):

k!